find two consecutive odd integers such that 47 more than the lesser is four times the greater.
Answer by Olasko(10)

Let the numbers be n and n+2 respectively.
47 + n = 4 ( n + 2 )
opening bracket
4n + 8 = 47 + n
collecting like terms
3n = 39
n = 13. The numbers are 13 and 15